Suppose that at 200 units of output a monopolist is producing such that marginal revenue is equal to marginal cost. The firm is selling its output at a price of $5 per unit and is incurring average variable costs of $3 per unit and average total costs of $4 per unit. On the basis of this information we can conclude that the firm:
is operating at maximum profit by producing the 200 units of output
should increase its use of variables inputs in order to reduce its total variable costs TVC
is operating at a loss that is less than the loss incurred by shutting down
should close down
1
Expert's answer
2014-12-09T09:56:48-0500
Q = 200 units at MR = MC, P = $5, AVC = $3 per unit, ATC = $4 per unit.
On the basis of this information total profit TP = (P - ATC)*Q = (5 - 4)*200 = $200, and as at the output of 200 units MR = MC, we can conclude that the firm is operating at maximum profit by producing the 200 units of output.
So, the right answer is a) is operating at maximum profit by producing the 200 units of output.